40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 22 16:15:52 2009, Chris Zimmerman, Update, General, Week 5/6 Update Beam_Scan.jpgLong_Range_Michelson_Setup_1_-_Actual.pngNSD_Displacement.png
    Reply  Sat Jul 25 13:34:58 2009, Koji, Update, General, Week 5/6 Update 
Message ID: 1779     Entry time: Wed Jul 22 16:15:52 2009     Reply to this: 1789
Author: Chris Zimmerman 
Type: Update 
Category: General 
Subject: Week 5/6 Update 

The last week I've started setting up the HeNe laser on the PSL table and doing some basic measurements (Beam waist, etc) with the beam scan, shown on the graph.  Today I moved a few steering mirrors that steve showed me from at table on the NW corner to the PSL table.  The goal setup is shown below, based on the UCSD setup.  Also, I found something that confused me in the EUCLID setup, a  pair of quarter wave plates in the arm of their interferometer, so I've been working out how they organized that to get the results that they did.  I also finished calculating the shot noise levels in the basic and UCSD models, and those are also shown below (at 633nm, 4mw) where the two phase-shifted elements (green/red) are the UCSD outputs, in quadrature (the legend is difficult to read).

 

 

Attachment 1: Beam_Scan.jpg  42 kB  | Show | Hide all | Show all
Attachment 2: Long_Range_Michelson_Setup_1_-_Actual.png  139 kB  | Hide | Hide all | Show all
Long_Range_Michelson_Setup_1_-_Actual.png
Attachment 3: NSD_Displacement.png  80 kB  | Show | Hide all | Show all
ELOG V3.1.3-